# A Grey-box Attack against Latent Diffusion Model-based Image Editing by Posterior Collapse

### v3

@@ Abstract
Recent advancements in generative AI, particularly Latent Diffusion Models (LDMs), have revolutionized image synthesis and manipulation. However, these generative techniques raises concerns about data misappropriation and intellectual property infringement. Adversarial attacks on machine learning models have been extensively studied, and a well-established body of research has extended these techniques as a benign metric to prevent the underlying misuse of generative AI. Current approaches to safeguarding images from manipulation by LDMs are limited by their reliance on model-specific knowledge and their inability to significantly degrade semantic quality of generated images. In response to these shortcomings, we propose the Posterior Collapse Attack (PCA) based on the observation that VAEs suffer from posterior collapse during training. Our method minimizes dependence on the white-box information of target models to get rid of the implicit reliance on model-specific knowledge. By accessing merely a small amount of LDM parameters, in specific merely the VAE encoder of LDMs, our method causes a substantial semantic collapse in generation quality, particularly in perceptual consistency, and demonstrates strong transferability across various model architectures. Experimental results show that PCA achieves superior perturbation effects on image generation of LDMs with lower runtime and VRAM. Our method outperforms existing techniques, offering a more robust and generalizable solution that is helpful in alleviating the socio-technical challenges posed by the rapidly evolving landscape of generative AI.
Figure 1: Comparison of our proposed method with AdvDM (Liang et al. 2023), PhotoGuard (Salman et al. 2023), MIST (Liang and Wu 2023), and SDS (Xue et al. 2024). Our method only requires access to the encoder of the VAE, achieving equal or superior performance on multi scenarios (Figure 2) with comparable image editing semantic degrade, lower runtime, and less occupied VRAM.
Figure 2: Comparison of our method with other baselines. The first column shows the original images to be manipulated. The top row displays the clean image and its manipulation results, while subsequent rows show attacked images and their corresponding outputs. We apply various prompts to simulate common scenarios: P1 (void text), P2 (BLIP-generated caption: “small toy astronauts standing next to each other” and “a raccoon dressed as lord”), P3 (weather change: “add some snow”), P4 (lighting adjustment: “apply sunset lighting”), and P5 (style transfer: “make it like a watercolor painting”). Our proposed method significantly disrupts image editing, causing substantial semantic loss. This is particularly evident in the results for P1, where the structural information of the edited image without prompt is completely lost.
Introduction
@@
The field of generative artificial intelligence has witnessed unprecedented advancements, particularly in the domain of image synthesis and manipulation. State-of-the-art methodologies, exemplified by diffusion models such as Stable Diffusion (Rombach et al. 2022), have demonstrated remarkable capabilities in producing photorealistic imagery with exceptional fidelity and diversity. These technological breakthroughs have not only revolutionized creative industries but have also democratized access to sophisticated image editing tools, empowering users across various domains.
@@
However, the proliferation of such powerful technologies inevitably engenders a concomitant set of ethical quandaries and potential security vulnerabilities. Of particular concern is the prospect of data misappropriation and intellectual property infringement. The facility with which diffusion-based models can be employed to manipulate existing visual content presents a significant challenge to the integrity of digital assets. For instance, malicious actors could exploit Stable Diffusion to edit copyrighted images, effectively “laundering” them by removing or altering identifying features (Rombach et al. 2022). This process of “de-copyrighting” not only undermines the rights of content creators but also poses a threat to the economic ecosystems built around digital imagery, potentially destabilizing industries ranging from photography to digital art.
@@
In the realm of machine learning security, adversarial attacks have emerged as a critical area of study (Lau et al. 2023a). These attacks aim to perturb the input of machine learning models in ways that are imperceptible to humans but can significantly disrupt the model’s output. Interestingly, this concept of adversarial attacks provides a potential solution to the misuse of generative AI. By introducing carefully crafted, acceptable perturbations to data that could be misused, it may be possible to impede the ability of diffusion models to edit or manipulate such content effectively.
@@
Current protection techniques often rely heavily on extensive prior knowledge, specifically requiring full white-box access to the parameters of the target models (Liang et al. 2023; Salman et al. 2023; Liang and Wu 2023; Xue et al. 2024). This assumption is largely impractical in real-world scenarios, especially given the rapid pace of technological advancement in generative AI. The proliferation of model architectures and constant iterations of backbone networks create a dynamic landscape where protection methods quickly become obsolete. Developing a universal method that can effectively protect against the vast majority of generative models without detailed knowledge of their internal structures is exceedingly difficult. Furthermore, the diversity of model architectures means that a method designed for one particular model or family may be entirely ineffective against others. This lack of generalizability severely limits the practical applicability of such approaches in a rapidly changing technological environment.
Secondly, while existing approaches successfully prevent infringer-desired alterations and introduce visible watermark-like artifacts in post-editing images as indicators of unauthorized manipulation (Liang et al. 2023; Xue et al. 2024), these approaches fail to comprehensively disrupt the semantic quality of generation as edited adversarial samples still remain the perceptual integrity with original clean image. Consequently, these methods struggle to achieve the goal of thoroughly degrading the semantic quality of manipulated outputs.
@@
Given these limitations in current protective measures and the narrow scope of existing attack methods on diffusion models, addressing these critical gaps in the field requires a solution that meets several key criteria:
@@
  • The method should cause significant degradation to the semantic generation quality of the most popular image editing and generation methods within the bounds of human visual sensitivity.
  • The method should exhibit strong transferability, easily generalizing across various models and architectures.
@@
In literature, existing attacks on VAEs primarily focus on corrupting the decoding process to lead to deteriorated reconstruction (Kuzina, Welling, and Tomczak 2022), which does not align with our objective of perturbing downstream diffusion process. Notably, the encoding posterior distribution of VAEs, often overlooked in previous attack methods, provides a more comprehensive representation of the latent space, which is the bottleneck of all downstream tasks, potentially leading to more effective and generalizable attacks.
Drawing inspiration from the phenomenon of posterior collapse in VAEs’ training (Razavi et al. 2019), where the latent variables fail to capture meaningful information about the input data, we suggest to intentionally induce this collapse as an attack strategy. Thus, we propose Posterior Collapse Attack (PCA), inducing collapse in the posterior distribution with our proposed new loss function, resulting worse semantic quality of edited image. As illustrated in Figure 1 and 2, our method effectively perturbs downstream LDMs in a near black-box manner (only 3.39% paprameters). This approach not only aligns with our goal of disrupting downstream tasks but also offers a more robust and transferable solution. Unlike previous methods that rely heavily on model-specific knowledge, our method merely leverages encoder of VAEs, providing a more universally applicable protection mechanism against unauthorized image manipulation across various generative AI architectures.
@@
  • We introduce the Posterior Collapse Attack that merely exploits the common VAE structure in LDMs, utilizing a novel loss function to cause the posterior collapse. Our method induces a significant semantic collapse of LDM-edited images, with minimal model-specific knowledge and less computational resources.
  • Our experimental results demonstrate that PCA, operating with near black-box adversarial information, effectively prevent LDM-based image editing across various architectures and resolutions, outperforming existing methods in terms of transferability and robustness.
Related Work
Generation Models
@@
Diffusion Probabilistic Model (DPM) (Ho, Jain, and Abbeel 2020) has achieved state-of-the-art results in density estimation (Kingma et al. 2021) and sample quality (Dhariwal and Nichol 2021). These models are powerful due to their U-Net backbone, which suits the inductive biases of image-like data. However, they face challenges in inference speed and training costs, especially for high-resolution images. To address these limitations, researchers have developed approaches like Denoising Diffusion Implicit Models (DDIM) (Song, Meng, and Ermon 2021) to enhance sampling speeds, and explored two-stage processes. Latent Diffusion Models (LDM) (Rombach et al. 2022) use autoencoding models to learn a perceptually similar space with lower computational complexity, while VQ-VAEs (Yan et al. 2021; Razavi, Van den Oord, and Vinyals 2019) and VQGANs (Yu et al. 2022; Esser, Rombach, and Ommer 2021) utilize discretized latent spaces and adversarial objectives to scale to larger images. These advancements aim to overcome the challenges of complex training, data shortages, and computational costs associated with diffusion models, particularly for high-resolution image synthesis.
Adversarial Attack
@@
The field of adversarial machine learning was catalyzed by the seminal work of Szegedy et al. (2014), who first uncovered the vulnerability of neural networks. Subsequent research has led to the development of various attack methodologies (Goodfellow, Shlens, and Szegedy 2015; Madry et al. 2018). The implications of adversarial attacks raised significant concerns in critical applications. This has spurred the development of defensive strategies, such as adversarial training  (Lau et al. 2023b) and input transformation (Lin et al. 2020b). We have seen the expansion of adversarial attack research into more complex domains (Liu et al. 2023; Guo et al. 2024a; Li et al. 2024; Zhao et al. 2024).
Attack on VAEs.
@@
Adversarial attacks on VAEs have attracted significant research attention. Most existing research on VAE attacks focused on manipulating the input or latent space to affect reconstruction output (Tabacof, Tavares, and Valle 2016; Gondim-Ribeiro, Tabacof, and Valle 2018; Kos, Fischer, and Song 2018; Kuzina, Welling, and Tomczak 2022), aiming to generate adversarial examples that would be reconstructed as images from different classes. Our approach, however, differs significantly from these previous works. We specifically target the disruption of downstream tasks based on VAE encodings, such as diffusion models, rather than the VAEs’ own reconstruction. Furthermore, we attack the parameters of the approximate posterior distribution produced by the encoder, instead of the latent variables themselves. This approach allows us to explore vulnerabilities in VAE-based systems, particularly in scenarios where VAEs serve as feature extractors or in applications relying on the distributional properties of encoded representations.
Attack on Diffusion-based Image Editing.
@@
Recent research has paid attention to adversarial attacks against diffusion models for image editing. Liang et al. (2023) targeted the noise prediction module, while Salman et al. (2023) and Shan et al. (2023) minimized latent space distances. Liang and Wu (2023) combined semantic and textural losses for protection. Xue et al. (2024) revealed the encoder’s vulnerability in latent diffusion models and proposed using Score Distillation Sampling to reduce computational costs. However, these approaches often rely heavily on full white-box access to target models or extensive knowledge of fundamental generative model principles. This dependency limits their practical applicability, especially given the rapid evolution of generative AI architectures. There is a need for methods that are less dependent on model-specific knowledge, can significantly degrade generation quality, and exhibit strong transferability across various models. Figure 1 shows the comparison of various methods across attacker’s knowledge of the generative model, runtime, and VRAM.
Preliminary
@@
Perceptual compression models typically use an autoencoder trained with a perceptual loss (Zhang et al. 2018) and a patch-based (Isola et al. 2017) adversarial objective (Dosovitskiy and Brox 2016). Specifically, given an image $\bm{x}\in\mathbb{R}^{H\times W\times 3}$ in pixel space, $\bm{x}$ is encoded into latent representation $z=\mathcal{E}(\bm{x})\in{\mathbb{R}}^{d}$ by an autoencoder. A decoder $\mathcal{D}$ reconstructs the image from the latent, giving $\tilde{\bm{x}}=\mathcal{D}(\bm{z})=\mathcal{D}(\mathcal{E}(\bm{x}))$, where $\bm{x}\in\mathbb{R}^{H\times W\times 3}$. The encoder downsamples the image by a factor $\mathit{f}=H/h=W/w$, $\mathit{f}$ is often set as 8 in latent diffusion models.
@@
$\bm{z}_{t}=\sqrt{\bar{\alpha}_{t}}\bm{z}_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon,\epsilon\sim\mathcal{N}(\mathbf{0},\mathbf{I}),$ (1)
where $\bar{\alpha}_{t}=\prod_{i=1}^{t}\alpha_{i},\alpha_{i}\in(0,1)$.
@@
$\mathbb{E}_{\bm{z}_{0},\bm{c},\epsilon,t}[\|\epsilon_{\theta}(\bm{z}_{t},\bm{c},t)-\epsilon\|],$ (2)
where $\epsilon_{\theta}$ is the model prediction. During inference, with a random Gaussian noise start $\bm{z}_{T}$ and condition $\bm{c}$, the diffusion model carries out the opposite process from step $T$ down to 1 to get the encoding of sampled image $\hat{\bm{z}}_{0}$ by:
$\hat{\bm{z}}_{t-1}=\frac{1}{\sqrt{\alpha_{t}}}\left(\hat{\bm{z}}_{t}-\frac{1-\alpha_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\epsilon_{\theta}(\hat{\bm{z}}_{t},\bm{c},t)\right)+\sigma_{t}\epsilon,$ (3)
where $\sigma_{t}=\frac{1-\bar{\alpha}_{t-1}}{1-\bar{\alpha}_{t}}\beta_{t},\beta_{t}=1-\alpha_{t}$.
Method
Problem Definition

### v4

# ruledlabelfont=normalfont,labelsep=colon,strut=off

@@ Abstract
Recent advancements in Latent Diffusion Models (LDMs) have revolutionized image synthesis and manipulation, raising significant concerns about data misappropriation and intellectual property infringement. While adversarial attacks have been extensively explored as a protective measure against such misuse of generative AI, current approaches are severely limited by their heavy reliance on model-specific knowledge and substantial computational costs. Drawing inspiration from the posterior collapse phenomenon observed in VAE training, we propose the Posterior Collapse Attack (PCA), a novel framework for protecting images from unauthorized manipulation. Through comprehensive theoretical analysis and empirical validation, we identify two distinct collapse phenomena during VAE inference: diffusion collapse and concentration collapse. Based on this discovery, we design a unified loss function that can flexibly achieve both types of collapse through parameter adjustment, each corresponding to different protection objectives in preventing image manipulation. Our method significantly reduces dependence on model-specific knowledge by requiring access to only the VAE encoder, which constitutes less than 4% of LDM parameters. Notably, PCA achieves prompt-invariant protection by operating on the VAE encoder before text conditioning occurs, eliminating the need for empty prompt optimization required by existing methods. This minimal requirement enables PCA to maintain adequate transferability across various VAE-based LDM architectures while effectively preventing unauthorized image editing. Extensive experiments show PCA outperforms existing techniques in protection effectiveness, computational efficiency (runtime and VRAM), and generalization across VAE-based LDM variants. Our code is available at https://github.com/ZhongliangGuo/PosteriorCollapseAttack.
I Introduction
@@
The field of generative artificial intelligence has witnessed unprecedented advancements, particularly in the domain of image synthesis and manipulation. State-of-the-art methodologies, exemplified by diffusion models such as Stable Diffusion [ldm], have demonstrated remarkable image generation quality and speed compared to other generation paradigm [li2024autoregressive]. These breakthroughs have revolutionized creative industries and democratized access to sophisticated image editing tools, empowering users across various domains.
@@
However, the proliferation of such powerful technologies inevitably engenders a concomitant set of ethical quandaries and potential security vulnerabilities. Of particular concern is the risk of data misappropriation and intellectual property infringement. The facility with which diffusion-based models can be employed to manipulate existing visual content presents a significant challenge to the integrity of digital assets. For instance, malicious actors could exploit Stable Diffusion to edit copyrighted images, effectively “laundering” them by removing or altering distinctive features [ldm]. This process of “de-copyrighting” not only undermines the rights of content creators but also poses a threat to the economic ecosystems built around digital imagery, potentially destabilizing industries ranging from photography to digital art.
@@
In the realm of machine learning security, adversarial attacks have emerged as a critical area of study [tifs1]. These attacks aim to perturb the input of models in ways that are imperceptible to humans but can significantly disrupt the output. This concept of adversarial attacks provides a promising solution to the misuse of generative AI [liu2025diffprotect]. By introducing carefully crafted, acceptable perturbations to data that could be misused, it may be possible to impede the ability of diffusion models to edit or manipulate such content effectively.
@@
Current protection techniques can be broadly categorized based on their objectives: some aim to preserve the original image by minimizing perturbation distance, while others maximize deviation from the expected edited output. However, these approaches often rely heavily on extensive prior knowledge, specifically requiring full white-box access to the parameters of the target models [advdm, photoguard, mist, sds]. Specifically, existing methods [advdm, photoguard, mist, sds] were naturally designed to target the U-Net, the primary generative component of LDMs, with their loss computations fundamentally requiring U-Net participation. While this focus on the core generation module is intuitive, it necessitates access much more model parameters and introduces substantial computational overhead. This assumption is largely impractical in real-world scenarios, especially given the rapid pace of technological advancement in generative AI. The proliferation of model architectures and continuous evolution of backbone networks create a dynamic landscape in which protection methods quickly become obsolete. Developing a universal method that can effectively protect against the vast majority of generative models without detailed knowledge of their internal structures is exceedingly difficult. Furthermore, the diversity of model architectures means that a method designed for one particular model or family may be entirely ineffective against others. This lack of generalizability severely limits the practical applicability of such approaches in a rapidly changing technological environment.
Figure 1: Comparison of our proposed method with AdvDM [advdm], PhotoGuard [photoguard], MIST [mist], and SDS [sds]. Our method only requires access to the encoder of the VAE, achieving equal or superior performance on multi scenarios (Fig. 3) with comparable image editing semantic degrade, lower runtime, and less occupied VRAM. All runtime and VRAM measurements were conducted on a NVIDIA RTX 3090 with $T=40$ iterations and 512×512 resolution.
Figure 2: Overview of the Posterior Collapse Attack (PCA). The VAE encoder maps input images to posterior distributions $q(\bm{z}|\bm{x})=\mathcal{N}(\bm{\mu},\bm{\sigma}^{2})$. By manipulating the KL divergence with different hyperparameter $v$, PCA induces two types of collapse: PCA+ (Objective 1) maximizes KL to create diffusion collapse, pushing the distribution out-of-distribution with $|\bm{\mu}|\to\infty$ and $\bm{\sigma}^{2}\to\infty$, thereby preserving original image content; PCA- (Objective 2) minimizes KL to create concentration collapse, collapsing the distribution to a Dirac $\delta$ function with $\bm{\mu}\to 0$ and $\bm{\sigma}^{2}\to 0$, thereby destroying semantic information.
@@
Given these limitations in current protective measures and the narrow scope of existing attack methods on diffusion models, addressing these critical gaps requires a solution that satisfies several key criteria:
@@
  • The protection mechanism should maintain high efficiency while requiring minimal computational overhead, making it practical for scenarios where computational resources may be constrained.
  • For the effectiveness, the method should exhibit better transferability, easily generalizing across various models and architectures.
@@
In literature, existing attacks on VAEs primarily focus on corrupting the decoding process, leading to deteriorated reconstruction [kuzina2022alleviating], which does not align with our objective of disrupting the downstream diffusion process. Notably, the encoding posterior distribution of VAEs, often overlooked in previous attack methods, provides a more comprehensive representation of the latent space, which is the bottleneck of all downstream tasks, potentially leading to more effective and generalizable attacks.
Drawing inspiration from the phenomenon of posterior collapse in VAEs’ training [razavi2018preventing], in which the latent variables fail to capture meaningful information from the input data, we propose the Posterior Collapse Attack (PCA) to intentionally induce this collapse-like phenomenon in the inference stage. As shown in Fig. 2, our exploration reveals two distinct types of collapse during inference: diffusion collapse and concentration collapse, each corresponding to a different protection objective (preserving image content versus disrupting edits). Through our proposed novel loss function, PCA can flexibly achieve either protection objective by inducing the corresponding type of collapse through parameter adjustment.
As illustrated in Fig. 2 and Fig. 3, our method effectively prevents unauthorized image manipulation in a near black-box manner (less than 4% parameters). This approach offers a more robust and transferable solution. Unlike previous methods that rely heavily on model-specific knowledge, our method merely leverages encoder of VAEs, providing a more universally applicable protection mechanism against unauthorized image manipulation across various generative AI architectures.
@@
  • We characterize two distinct collapse phenomena in VAE inference, diffusion collapse and concentration collapse, establishing their connection to protection objectives and bridging collapse mechanisms with protection strategies.
  • We design a unified loss function achieving both protection types through parameter adjustment, eliminating separate methods for different goals.
  • Extensive experiments demonstrate that PCA achieves strong protection performance while requiring minimal white-box information, and exhibits better transferability across various diffusion models and architectures.
II Related Work
II-A Generation Models
@@
Diffusion Probabilistic Model (DPM) [ho2020ddpm] has achieved state-of-the-art results in density estimation [kingma2021variational] and sample quality [dhariwal2021diffusionbeatgan]. These models are powerful due to their U-Net backbone, which suits the inductive biases of image-like data. However, they face challenges in inference speed and training costs, especially for high-resolution images. To address these limitations, researchers have developed approaches like Denoising Diffusion Implicit Models (DDIM) [song2020ddim] to enhance sampling speeds, and explored two-stage processes. Latent Diffusion Models (LDM) [ldm] use autoencoding models to learn a perceptually similar space with lower computational complexity, while VQ-VAEs [yan2021videogpt, razavi2019vqvae2] and VQGANs [yu2021vectorvqgan, vqgan] utilize discretized latent spaces and adversarial objectives to scale to larger images. These advancements aim to overcome the challenges of complex training, data shortages, and computational costs associated with diffusion models, particularly for high-resolution image synthesis.
II-B Adversarial Attack
@@
The field of adversarial machine learning was catalyzed by the seminal work of Szegedy et al. [szegedy2013intriguing], who first uncovered the vulnerability of neural networks. Subsequent research has led to the development of various attack and defense methodologies [goodfellow2014explaining, madry2017towards, tifs3]. The implications of adversarial attacks raised significant concerns in critical applications. This has spurred the development of defensive strategies, such as adversarial training  [lau2023interpolated, tifs2] and input transformation [lin2020dual]. We have seen the expansion of adversarial attack research into more complex domains [liu2023instruct2attack, guo2024white, li2024threats, zhao2024survey].
II-B1 Attack on VAEs
@@
Early works of adversarial attacks on VAEs primarily focused on attacking the image reconstruction capabilities of VAEs. Various attack methods against VAE and VAE-GAN models were developed [tabacof2016adversarial, gondim2018adversarial, kos2018adversarial], including classifier-based attacks, attacks using the VAE loss function, and latent space attacks. These methods aimed to generate adversarial examples that would be reconstructed as images from different classes. Recent research has expanded to investigating attacks on the encoder output of hierarchical VAEs, considering the higher-level latent variables [kuzina2022alleviating]. Most existing research on VAE attacks has focused on manipulating the input or latent space to affect reconstruction output. Our approach, however, differs significantly from these previous works. We specifically target the disruption of downstream tasks based on VAE encodings, such as diffusion models, rather than the VAEs’ own reconstruction. Furthermore, we attack the distributional parameters ($\bm{\mu},\bm{\sigma}^{2}$) defining the approximate posterior $q(\bm{z}|\bm{x})$, rather than the sampled latent variables themselves. This allows us to explore vulnerabilities in VAE-based systems where distributional properties are critical for downstream processing.
II-B2 Attack on Diffusion-based Image Editing
@@
Recent research has paid attention to adversarial attacks against diffusion models for image editing. [advdm] targeted the noise prediction module, while [photoguard] and [shan2023glaze] minimized latent space distances. [mist] combined semantic and textural losses for protection. [sds] revealed the encoder’s vulnerability in latent diffusion models and proposed using Score Distillation Sampling to reduce computational costs. However, as shown in Fig. 2, these approaches often rely heavily on full white-box access to target models or extensive knowledge of fundamental generative model principles. Specifically, the design of existing methods fundamentally requiring U-Net participation alongside Text Encoder and (optionally) VAE Decoder. They exploit behavioral vulnerabilities in generation process, which cannot be trivially adapted to methods using only VAE encoder that targets architectural vulnerabilities in posterior distribution. This dependency limits their practical applicability, especially given the rapid evolution of generative AI architectures. There is a need for methods that are less dependent on model-specific knowledge, can significantly degrade generation quality, and exhibit better transferability across various models.
Figure 3: Comparison of our method with other baselines subject to different Objectives. The first column shows editing prompts. The second column shows the original input images and expected edited outputs. The subsequent columns show images protected by different methods and their corresponding outputs. Notably, $x$ refers to the original image will be edited by LDM; $f(\cdot)$ refers to the LDM-based image editing.
III Preliminary
@@
Perceptual compression models typically use an autoencoder trained with a perceptual loss [zhang2018unreasonable] and a patch-based [isola2017image] adversarial objective [dosovitskiy2016generating]. Specifically, given an image $\bm{x}\in\mathbb{R}^{H\times W\times 3}$ in pixel space, $\bm{x}$ is encoded into latent representation $z=\mathcal{E}(\bm{x})\in\mathbb{R}^{d}$ by an autoencoder. A decoder $\mathcal{D}$ reconstructs the image from the latent, giving $\tilde{\bm{x}}=\mathcal{D}(\bm{z})=\mathcal{D}(\mathcal{E}(\bm{x}))$, where $\bm{x}\in\mathbb{R}^{H\times W\times 3}$. The encoder downsamples the image by a factor $\mathit{f}=H/h=W/w$, $\mathit{f}$ is often set as 8 in latent diffusion models.
@@
$\bm{z}_{t_{\text{diff}}}=\sqrt{\bar{\alpha}_{t_{\text{diff}}}}\bm{z}_{0}+\sqrt{1-\bar{\alpha}_{t_{\text{diff}}}}\epsilon,\epsilon\sim\mathcal{N}(\mathbf{0},\mathbf{I}),$ (1)
where $\bar{\alpha}_{t_{\text{diff}}}=\prod_{i=1}^{t_{\text{diff}}}\alpha_{i},\alpha_{i}\in(0,1)$.
@@
$\mathbb{E}_{\bm{z}_{0},\bm{c},\epsilon,{t_{\text{diff}}}}[\|\epsilon_{\theta}(\bm{z}_{t_{\text{diff}}},\bm{c},{t_{\text{diff}}})-\epsilon\|],$ (2)
where $\epsilon_{\theta}$ is the model prediction. During inference, with a random Gaussian noise start $\bm{z}_{T_{\text{diff}}}$ and condition $\bm{c}$, the diffusion model carries out the opposite process from step $T_{\text{diff}}$} down to 1 to get the encoding of sampled image $\hat{\bm{z}}_{0}$ by:
$\hat{\bm{z}}_{{t_{\text{diff}}}-1}=\frac{1}{\sqrt{\alpha_{t_{\text{diff}}}}}\left(\hat{\bm{z}}_{t_{\text{diff}}}-\frac{1-\alpha_{t_{\text{diff}}}}{\sqrt{1-\bar{\alpha}_{t_{\text{diff}}}}}\epsilon_{\theta}(\hat{\bm{z}}_{t_{\text{diff}}},\bm{c},{t_{\text{diff}}})\right)+\sigma_{t_{\text{diff}}}\epsilon,$ (3)
where $\sigma_{{t_{\text{diff}}}}=\frac{1-\bar{\alpha}_{{t_{\text{diff}}}-1}}{1-\bar{\alpha}_{{t_{\text{diff}}}}}\beta_{{t_{\text{diff}}}},\beta_{{t_{\text{diff}}}}=1-\alpha_{{t_{\text{diff}}}}$.
IV Method
IV-A Problem Definition